solve this system of equations 4a+7b=-30 8a+4c=-40 6b+4c=-20
Answer by TimothyLamb(4379)

4a + 7b = -30
8a + 4c = -40
6b + 4c = -20
---
change to standard variable names x y z
---
4x + 7y = -30
8x + 4z = -40
6y + 4z = -20
---
copy and paste the above standard form equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x=a= -4
y=b= -2
z=c= -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php